Wednesday, December 11, 2019, 11:00 EST
From Hall A Wiki
*********BLUEJEANS CALL-IN INSTRUCTIONS********* ---------------------------------------------------- 1. Dial Toll-Free Number: 1-888-240-2560 (U.S. & Canada) 2. International participants dial: Toll Number: 1-408-740-7256 3. Enter meeting ID:140287970 Desktop sharing: https://bluejeans.com/140287970
Meeting room = Cebaf Center F227
- Need to pick new meeting time
- Just received survey report for laser table
- Detector status
- QWAD adapter boards
- Counting DAQ
- Cavity/laser status
- CREX preparations
- Too busy for analysis!
Cip, Dave, Mark, Adam, Alex, Don, Jinlong, JC, Iris
- Meeting moved to 10 am
- Offset for the BPMs found. Worst case 1 mm. BPM offsets will be uploaded, Dave to take care of this.
- Collimator was checked in October, wait for survey data.
- Cavity is not locking. Mode matching is off.
- Changed the fibre amp (tried different see laser but that didn't improve things.)
- Today realigning PPLN to get a better shape. (but highest priority is to get to high current.)
- Electron Detector
- FSD is fixed
- QWAD adaptor tested and seems to work. Will insert when beam is threaded.
- Progress made on scaler readout butt still issues to resolve.
- DAQ is generally is good shape.
- Is helicity signal in the DAQ?
- Need to work on the software (decoding of VTROC.)
- Photon Detector
- Looking into a waterproofing solution. Current plastic doesn't fit, try to get a bigger piece of plastic.
- All cable ends have been replaced.
- Finger scintillators behind each tungsten finger was checked (take a picture.)
- Updating online analysis for CREX (cosmetic)
- Make rootfiles available for use in PREX background studies.